Show an equation and a solution. How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.)
Answer by edjones(8007)

12 * .4 = 4.8
x * 1.0 = x
(x+12)*.6=4.8x
.
.6x+7.2=4.8x
-4.2x=-7.2
x=7.2/4.2
=12/7 L
.
Ed